OpenStudy (anonymous):

a triangle has sides of 3, 6, and 8. In a similar triangle the shortest side is 9 and the longest side is x. a) write a proportion that models the situation. b) solve the proportion for x. need a little guidance on this one please...

OpenStudy (inkyvoyd):

3:6:8 9:a:x

OpenStudy (inkyvoyd):

Multiply the known ratio by three to get the side length values for the unknown triangle.

OpenStudy (inkyvoyd):

If you're having trouble grasping what I'm saying feel free to ask questions for clarification

OpenStudy (anonymous):

no i'm trying to work it out first on paper, then i'll ask for a check :) lol

OpenStudy (anonymous):

so far I have k=-13/2 working on the rest... lol for y=

OpenStudy (inkyvoyd):

otay - tell me your final answer so I can check it :)

OpenStudy (anonymous):

which I got -39

OpenStudy (anonymous):

so k= -13/2 and y=-39 .... hopefully LOL

OpenStudy (anonymous):

@inkyvoyd

OpenStudy (inkyvoyd):

well, that side length you should be getting is 27 - it's a scale of 3...

OpenStudy (anonymous):

oh...sorry...I was posting the answer to my next question... my bad.

OpenStudy (anonymous):

got ahead of myself lol

OpenStudy (inkyvoyd):

ohh lol

OpenStudy (anonymous):

ok..... on this one I got 3/8=9x and then came up with the longest side being 24. @inkyvoyd yes? LOL

OpenStudy (inkyvoyd):

uhh, you should be getting 27 I think :S

OpenStudy (anonymous):

i submitted my answer I got it right. lol thanks for the help :)

OpenStudy (inkyvoyd):

lol, then I epically fail. My bad.

OpenStudy (anonymous):

it's ok :)